sotjo 60-jyou 100tjoon 220 Vrms bo H2 Find Apparent power and Power factor.
Answer #1

solution 220 voms Z₃ = 60-404 Z2 = 100+601 23= 50 Power V. Ik I= V Zeq. Zeq = Zil/22/23 zi - Z1 Z2 ZitZ2 = 160-aj (100+600) 1zeqa zi.23 ZitZz - (5138-8.924) 50 101-38-8.923 Zeq= 25.6203 /-4.82° Now, Power V It V. Vt Zeq - Vouto 25.6203 +482 . V=220 Lhere we calculated equalent impedance of parallel circuit. And then calculated apparent power and power factor.